Grade 12Mechanics

The rate of maximum velocity and maximum displacement of a particle executing simple harmonic motion is equal to

SNAPDEEP
6 Years agoGrade 12

1 Answer

Arun
6 Years ago
 
yes, \frac{v_{max}}{x_{max}}=\omega  .U can derive the formula v=\omega \sqrt{A^{2}-x^{2}} by phasor diagram easily.
clearly for v to be max. x^{2}​ has to be min. so, v is max. when x=0. Hence, v_{max}=\omega A.
and x_{max} is obviosly equal to A. Hence,\frac{v_{max}}{x_{max}}=\omega